OpenStudy (bee_see):

Compute (4 + 2^200) mod 3.

jimthompson5910 (jim_thompson5910):

2^(200) = 2^(2*100) (mod 3) 2^(200) = (2^2)^100 (mod 3) 2^(200) = 4^100 (mod 3) 2^(200) = 1^100 (mod 3) ... since 4 = 1 (mod 3) 2^(200) = 1 (mod 3)

OpenStudy (bee_see):

why do you want to get 2^(2*100) (mod 3)?

jimthompson5910 (jim_thompson5910):

because I wanted to get to (2^2)^100

jimthompson5910 (jim_thompson5910):

I used the identity (a^x)^y = a^(x*y)

OpenStudy (bee_see):

My professor has an example so I set it up like that and I got stuck on (1 mod 3 + 2^200) mod 3.

OpenStudy (bee_see):

why would you want to get (2^2)^100?

jimthompson5910 (jim_thompson5910):

because I know that 4 = 1 (mod 3)

jimthompson5910 (jim_thompson5910):

so 4^100 = 1^100 (mod 3) 1 to any power = 1, so 4^100 = 1^100 = 1 (mod 3)

OpenStudy (bee_see):

how is 4 = 1 (mod 3)?

jimthompson5910 (jim_thompson5910):

because they both leave the same remainder when divided by 3 4/3 = 1 remainder 1 1/3 = 0 remainder 1

jimthompson5910 (jim_thompson5910):

another example 10 = 4 (mod 3) since 10/3 = 3 remainder 1 4/3 = 1 remainder 1 or you can use this rule a = b (mod m) if and only if m | (a-b) in english: if m is a factor of (a-b) then a = b (mod m)

jimthompson5910 (jim_thompson5910):

using this rule `a = b (mod m) if and only if m | (a-b)` we know 4 = 1 (mod 3) because 3 is a factor of (4-1) same with 10 = 4 (mod 3) since 3 is a factor of (10-4)

OpenStudy (bee_see):

how did 2^(2*100) (mod 3 go to (2^2)^100 (mod 3)? I don't really get what you are trying to obtain at the end.

jimthompson5910 (jim_thompson5910):

Like I said, I used this identity `(a^x)^y = a^(x*y)`

ganeshie8 (ganeshie8):

its not much different, but u may also work it like this : (4 + 2^200) mod 3 = `4 mod3` + `2^200 mod3` = `1` + `(-1)^200 mod3` = `1` + `1` = 2

jimthompson5910 (jim_thompson5910):

@ganeshie8 's method is much more effective

OpenStudy (bee_see):

do did you get (-1)^200 mod3?

OpenStudy (bee_see):

how did*

jimthompson5910 (jim_thompson5910):

2 = -1 (mod 3) since they both lie in the same equivalence class

jimthompson5910 (jim_thompson5910):

set of numbers that when you divide them by 3, you get remainder 0 {..., -9, -6, -3, 0, 3, 6, 9, ...} set of numbers that when you divide them by 3, you get remainder 1 {..., -8, -5, -2, 1, 4, 7, 10, ...} set of numbers that when you divide them by 3, you get remainder 2 {..., -7, -4, -1, 2, 5, 8, 11, ...} these sets are known as equivalence classes so for example, 4 and 7 are from the same equivalence class mod 3, which is why 4 = 7 (mod 3)

ganeshie8 (ganeshie8):